OpenStudy (kylie.diprimo):

Have absolutely no idea how to do this A circle’s area is represented by A = π(x^2 − 22x + 121). What expression represents the radius of the circle? A. x + 11 B. x − 11 C. x − 22 D. x + 22 MEDAL!!!

OpenStudy (abdullahm):

\(\Large\bf Area = \pi r^2\) \(\Large\bf (a - b)^2 = a^2 - 2ab + b\) \(\Large\bf x^2 - 22x + 121 = ?\)

OpenStudy (abdullahm):

Whatever you factor it down to is the radius!

OpenStudy (kylie.diprimo):

I probably sound really stupid, but I don't get it

OpenStudy (abdullahm):

No, it's ok! Do you know how to factor a trinomial?

OpenStudy (kylie.diprimo):

Yes

OpenStudy (abdullahm):

All you need to do is factor x^2 - 22x + 121

OpenStudy (abdullahm):

Hint: -11 and -11 add up to -22 -11 and -11 multiply to 121

OpenStudy (kylie.diprimo):

Is it A)x+11

satellite73 (satellite73):

no because \((x+11)^2=x^2+22x+121\)

OpenStudy (kylie.diprimo):

Duh, D) x-22 right

OpenStudy (abdullahm):

No, \(\ (x-22)^2 = x^2 - 44x + 484\)

OpenStudy (kylie.diprimo):

Oy vey

OpenStudy (kylie.diprimo):

Is it B

OpenStudy (abdullahm):

Yes

OpenStudy (kylie.diprimo):

Thank you I am so embarrassed at my stupidity